OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (1 point) 2x + 9y = 4 3x + 7y = –7

OpenStudy (anonymous):

http://www.youtube.com/watch?v=0P0SCQf-hWQ

OpenStudy (anonymous):

check that out!

OpenStudy (campbell_st):

multiply equation (1) by 3 (6x + 27y = 12) multiply equation (2) by 2 (6x +14y = -14) the subtract (1) - (2) (13y = 26) divide both sides by 13 (y = 2) substitute into (1) (3x + 18 = 4) subtract 18 from both sides (3x = -14) divde both sides by 3 (x = -14/3)

OpenStudy (campbell_st):

oops error with finding x 2x + 18 = 4 2x = -14 x = -7